OpenStudy (anonymous):

Solve x2 − 4x = 12. x = −3, x = 4 x = 3, x = −4 x = 6, x = −2 x = −6, x = 2

OpenStudy (anonymous):

I'm not to sure about this.

OpenStudy (shamil98):

x^2 - 4x = 12 x^2 - 4x - 12 = 0 Factor it and then solve for x.

OpenStudy (anonymous):

Can you help me factor it

OpenStudy (shamil98):

What values when multiplied give you -12 and when added give you -4.

OpenStudy (anonymous):

-2?

OpenStudy (shamil98):

x*y = -12 x + y = -4 you could think of it as this.. or you could look it like this 6* -2 = -12 6 + (-2) = -4 (x+6)(x-2) = 0 x = ? , x = ?

OpenStudy (anonymous):

so, C? @shamil98